### v1

@@ 5 Research Question Results
Pre-trained SafeCoder Secure-Instruct Pre-trained SafeCoder Secure-Instruct
Secure (Vul, Secure) Secure (Vul, Secure)
@@
RQ2: How is Secure-Instruct compared with the state-of-the-art (SOTA) approach in securing LLM-based code generation?
Motivation. SafeCoder (He et al., 2024) represents the state-of-the-art in secure LLM-based code generation. It employs an automated pipeline to mine vulnerable code and corresponding fixes from a large corpus of open-source commits (145 million), followed by human intervention for post-processing and validation. The final dataset consists of (instruction, vulnerable code, secure code) triples used to fine-tune LLMs. In contrast, our proposed approach, Secure-Instruct , is a fully automated technique with zero human intervention. Given these fundamental differences, it is important to compare the performance of Secure-Instruct against SafeCoder to assess the effectiveness and scalability of our method.
Approach. We further conducted a comparative analysis between Secure-Instructsecure and SafeCoder on CWEBench and CWEval. In particular, we used the two released models, i.e., CodeLlma-7B and Mistral-7B, which were already fine-tuned by SafeCoder. In CWEBench, there are 44 CWE types. However, the training dataset of SafeCoder does not include all of them. SafeCoder was fine-tuned on a dataset containing 23 CWEs, whereas Secure-Instruct was fine-tuned on a dataset covering 44 CWEs. There are 20 CWEs covered by both Secure-Instruct and SafeCoder. We used the same configuration as described in the SafeCoder paper, namely a temperature of 0.8 and 100 samples for each security-specific scenario.
Results. Table 3 shows that Secure-Instruct achieves higher Func-Sec@1 on CWEval than SafeCoder for both CodeLlama‑7B and Mistral‑7B. Notably, SafeCoder exhibits a performance drop on CWEval. For example, SafeCoder’s CodeLlama-7B achieves only 7.1% Func-Sec@1, which is lower than the pre-trained model. A similar decline is observed in SafeCoder’s Mistral-7B. These findings are consistent with the observations reported in the study Peng et al. (2025).
Table 5 breaks down the detailed results on CWEBench of CodeLlama-7B and Mistral-7B in their pre-trained, SafeCoder, and Secure-Instructsecure, on each CWE in our evaluation benchmark. The highest secure ratio for each CWE is highlighted in bold. In the scenarios whose CWEs are covered in both approaches (the top section of the table), our analysis shows that CodeLlama-7B fine-tuned with Secure-Instructsecure achieved the highest overall secure ratio (70.6%), outperforming SafeCoder by 2.3% in secure ratio, Mistral-7B fine-tuned with Secure-Instructsecure achieved 68.9% secure ratio, with a 3.7% higher than SafeCoder. In the generalization evaluation (i.e., with the full security-specific scenarios of our benchmark), Secure-Instructsecure demonstrates strong generalization capabilities across a broader range of CWEs for both CodeLlama-7B and Mistral-7B. Specifically, CodeLlama-7B with Secure-Instructsecure significantly outperforms SafeCoder, achieving a 9.2% improvement, while Mistral-7B achieves a 6.0% increase. At the CWE category level, CodeLlama‑7B fine‑tuned with Secure-Instructsecure achieves the highest or ties for highest secure ratio in 29 of 44 CWEs, outperforming both its pretrained baseline and the SafeCoder‑fine‑tuned variant. Similarly, Mistral‑7B with Secure-Instructsecure leads in 25 of 44 CWEs relative to its pretrained and SafeCoder‑fine‑tuned counterpart. These results demonstrate that Secure-Instructsecure significantly enhances secure code generation, outperforming the SOTA approach, SafeCoder.
To assess statistical significance, we conducted Fisher’s exact tests comparing the number of secure code snippets generated by the pre-trained model and the model fine-tuned with Secure-Instructsecure for each CWE. CWEs with statistical significance (p<<0.05) are highlighted in bold, such as CWE-089. Notably, more than half of the CWEs in Secure-Instructsecure exhibit statistical significance, further demonstrating the robustness of our approach in enhancing secure code generation.
Summary of RQ2: Safecoder encounters a degradation in CWEval benchmark. Secure-Instructsecure significantly outperforms SafeCoder with 9.2% and 6.0% improvements across CWEBench. Even for those CWEs that are part of SafeCoder’s fine-tuning dataset, Secure-Instructsecure achieves 2.3% and 3.7% higher secure ratios in CodeLlama-7B and Mistral-7B, respectively.
RQ3: Can the LLM fine-tuned by Secure-Instruct preserve the functional correctness in code generation?
Motivation. A critical concern in fine-tuning LLMs in secure code generation is understanding how fine-tuning impacts the functionality of the generated code. While fine-tuning with secure code examples aims to improve the security aspects of the output, it may compromise the model’s ability to produce functionally correct code.
Approach. The benchmark, CWEval, evaluates the functionality of generated code by test cases with the metric Func@1. Besides, we evaluate the fine-tuned models on HumanEval and MBPP, two widely used benchmarks for assessing code generation in Python. For each task of HumanEval and MBPP, we generate 40 samples using both pre-trained and fine-tuned models and assess their functionality using the Pass@k metric, including Pass@1 and Pass@10. Pass@k is a standard metric for evaluating code generation, which represents the probability that at least one of the $k$ generated candidates for a given problem is correct.
Results. Table 3 presents the Func@1 results for all models. All models, except for StarCoder-1B, have improvements after using Secure-Instruct. CodeLlama-7B has an over 10% increase in Func@1 compared with the pre-trained model. Table 6 presents the Pass@1 (Pass@10 is available in our replication package) on the HumanEval and MBPP benchmarks for pre-trained LLMs, fine-tuned LLMs by SafeCoder and Secure-Instruct. For CodeLlama‑7B, among the pretrained, SafeCoder‑fine‑tuned, Secure-Instruct(vul, secure)‑fine‑tuned, and Secure-Instructsecure‑fine‑tuned variants, Secure-Instructsecure-fine-tuned model achieves the highest Pass@1 on HumanEval at 35.6%. For Mistral‑7B, the Secure-Instruct(vul, secure)‑fine‑tuned model achieves the highest score on MBPP at 30.7%. Overall, most models show performance improvements on both datasets, except for StarCoder-1B. On CWEval, its Func@1 dropped from 15.3% to 13.8% with Secure-Instruct(vul, secure), and 12.7% with Secure-Instructsecure. On HumanEval, the Pass@1 score of StarCoder-1B dropped from 11.7% to 10.9% with Secure-Instructsecureand from 11.7% to 6.7% with Secure-Instruct(vul, secure). This decline may be due to the smaller size of StarCoder-1B compared to the other LLMs.
Summary of RQ3: Secure-Instruct enhances secure code generation without sacrificing functional correctness. Across diverse LLMs and benchmarks, Secure-Instruct consistently outperforms both pre-trained LLMs and those fine-tuned by SafeCoder in generating functionally correct code.
Model Benchmark Pre-trained SafeCoder Secure-Instruct
(vul,secure) secure
CodeLlama- 7B HumanEval 25.5% 32.0% 34.7% 35.6%
MBPP 28.3% 33.7% 36.0% 34.7%
CodeGen2- 7B HumanEval 15.3% NA 8.1% 16.3%
MBPP 20.4% NA 15.4% 22.8%
Mistral- 7B HumanEval 23.0% 31.3% 31.6% 25.9%
MBPP 24.7% 29.3% 30.7% 27.1%
StarCoder- 1B HumanEval 11.7% NA 6.7% 10.9%
MBPP 13.7% NA 12.2% 17.7%
Table 6: Pass@1 results on HumanEval and MBPP for the pre-trained LLMs, SafeCoder fine-tuned LLMs, and Secure-Instruct (ours) fine-tuned LLMs.
6 Discussions
The accuracy of CodeQL. In CWEBench, the ground-truth of security is based on CodeQL, which is a static tool and may produce false positives. To address this concern, we randomly sampled 291 code snippets labeled as secure by CodeQL from four models and manually verified their security. All sampled snippets were confirmed to be truly secure. Additionally, we used another benchmark, CWEval, which evaluates security through test cases, to further validate our results.
The similarity between Secure-Instruct’s fine-tuning datasets and our evaluation benchmark. To clarify data leaks (i.e., overlap between fine-tuning dataset and benchmark), we performed a similarity analysis. Specifically, we first extracted all the secure code examples from the CodeQL repository for the 44 CWEs of our benchmark. Since most scenarios of the 44 CWEs are created based on the CodeQL repository, such secure code can be considered as ground-truth code for conducting such a similarity analysis. We compare each LLM-generated secure code with the secure code examples from the CodeQL repository under the same CWE. We then calculated the cosine similarity of each pair using TF-IDF embeddings (Sparck Jones, 1972). As there exist multiple secure examples for one CWE in the CodeQL repository, we selected the highest score, as the similarity score for an LLM-synthesized secure code snippet.
Figure 6 depicts the average cosine similarity score for the data generated by Secure-Instructsecure is 0.17, while the average score for data generated by Secure-Instruct(vul, secure) is 0.23. Secure-Instructsecure’s values align with findings from previous work (Wei et al., 2024) on synthesizing fine-tuning data using LLMs, which reported similarity scores typically ranging between 0.1 and 0.2 in the previous synthesis techniques. In contrast, Secure-Instruct(vul, secure) exhibits a slightly elevated similarity score, exceeding this range. These results indicate that Secure-Instructsecure generates data with lower similarity compared to Secure-Instruct(vul, secure). Importantly, this result indicates that the improvements from Secure-Instruct are not simply due to the inclusion of data from the same distribution of the CodeQL repository. Instead, the generated data provides additional diversity for fine-tuning the dataset.
Figure 4: Cosine similarities between Secure-Instruct’s fine-tuning dataset and the secure code examples by CodeQL (i.e., our evaluation benchmark).
Low cost of Secure-Instruct in generating fine-tuned datasets. We show that Secure-Instruct is able to generate high-quality fine-tuning data with low monetary cost and no human intervention. For example, when using ChatGPT (GPT-4o-2024-08-06), Secure-Instruct costs \$114.09 USD for generating all secure code snippets. Similarly, for Claude, the cost is \$109.74 USD for generating all secure code snippets. Compared to manually curated datasets, this synthesis method is highly cost-efficient. The cost per CWE-language is \$2.87 USD. Compared to our Secure-Instruct, prior work SafeCoder (He et al., 2024) imposes challenges on a high-efficient data curation process, i.e., a sharp reduction from hundreds of millions of code commits on open source repositories on GitHub to a total of 25,000 repositories that can be compiled and scanned by CodeQL for automatically detecting vulnerability-fixing commits, finally to merely 1,200 pairs of vulnerable and secure code examples. In addition, SafeCoder relies on manual effort to identify 465 high-quality code samples from the 1,211 samples produced by the automated pipeline. In contrast, Secure-Instruct demonstrates significant advancements in efficiency by reducing the time and cost required to construct high-quality datasets while maintaining a strong focus on generating reliable and secure code samples.
CodeLlama- 7B CodeGen2- 7B Mistral- 7B StarCoder- 1B
Secure-Instruct(vul, secure) 64.4% 49.0% 61.3% 63.4%
Secure-Instructsecure 61.9% 49.1% 62.0% 66.2%
7 Threats to Validity
@@
Internal Threats. LLMs can produce different outputs for the same prompt. To prevent the huge variations, we set a large sample size to ensure the stabilization results. CWEBench relies on CodeQL to detect vulnerability. Such a static analyzer may produce false positives. Although we manually reviewed samples, misclassifications could still influence the conclusions. We performed de-duplication across both our fine-tuning dataset and evaluation benchmarks (CWEBench and CWEval). However, this process may fail to detect duplicates, leading to false negatives.
@@
External Threats. We follow prior work and choose LLMs with smaller parameter sizes to make fine-tuning feasible under limited computational resources. Fine-tuning larger models, such as ChatGPT-4o, would incur significantly higher costs. As a result, we acknowledge that our findings may not fully generalize to larger LLMs, and we consider this a direction for future investigation. We include 44 distinct CWEs in this study, reflecting a broad range of vulnerabilities. However, numerous additional CWEs remain unaddressed in this study. Moreover, individual CWE can manifest in various real-world scenarios, further highlighting the complexity of comprehensive evaluation. To improve coverage, we expanded our security-sensitive prompts to incorporate additional CWEs beyond those included in earlier studies. Compared to prior work (Pearce et al., 2022; He and Vechev, 2023; He et al., 2024; Li et al., 2024a), our study evaluates a significantly greater number of CWEs, providing a more thorough assessment of generated code vulnerabilities. Future efforts should include broader CWEs. Additionally, we include six programming languages in our evaluation benchmark. While these languages are highly popular and widely adopted in software development, our findings may not generalize to all programming languages. Future work should expand the language set to strengthen the robustness of the results.
8 Conclusion
In this paper, we proposed Secure-Instruct, a novel framework to improve LLM for secure code generation by harnessing LLMs for generating instruction tuning datasets. Secure-Instruct overcomes the limitations of the state-of-the-art technique, particularly addressing dataset imbalance, inefficient data curation, and the significant manual effort required. Unlike existing approaches, Secure-Instruct can be easily adapted to support new CWEs. Additionally, we contribute CWEBench, the largest evaluation benchmark to date for secure code generation. Our evaluation shows that Secure-Instruct significantly improves the performance of four representative LLMs, achieving 14.3% more secure code on average on CWEBench and +5.3% Func-Sec@1 on average on CWEval. Compared to the state-of-the-art, Secure-Instruct demonstrates strong performance even with a smaller fine-tuning dataset, highlighting its scalability and efficiency. Finally, we show that LLMs fine-tuned with Secure-Instruct maintain comparable or improved functional correctness in generated code.
Statements and Declarations
Funding. This work was supported by the Fonds de recherche du Québec (Grant No. 2024-NOVA-346499) [^1] https://repertoire.frq.gouv.qc.ca/offres/detailOffres.do;jsessionid=C333738136C878333C20227CC5EB4E1B?methode=afficher, the Natural Sciences and Engineering Research Council of Canada (NSERC) Discovery Grant (RGPIN-2019-07007), and the NSERC CREATE Grant (555406-2021). We gratefully acknowledge their support.
Competing Interests. The authors have no financial or proprietary interests in any material discussed in this article.
Data, Materials, and/or Code Availability. The data and code used in this study are available at https://zenodo.org/records/15019488.
Authors’ Contributions. Junjie Li - Conceptualization, Methodology, Data collection, Analysis, Writing original draft. Fazle Rabbi - Material preparation, Data collection, Analysis. Bo Yang - Material preparation, Data collection, Analysis, Writing original draft. Jinqiu Yang - Conceptualization, Methodology, Writing review & editing, Supervision, Supervising experiment design and execution. Song Wang - Conceptualization, Methodology, Writing review & editing, Supervision, Supervising experiment design and execution.

### v2

@@ 5 Discussions
6 Conclusion
We presented Secure-Instruct, a novel framework for improving secure code generation by leveraging LLMs to synthesize instruction-tuning datasets. Secure-Instruct overcomes the key limitations of the SOTA technique, particularly addressing dataset imbalance, inefficient data curation, and the significant manual effort required. Unlike existing approaches, Secure-Instruct can be easily adapted to support new CWEs. We also introduce CWEBench, the largest benchmark for secure code generation. Our evaluation shows that Secure-Instruct significantly improves the performance of four LLMs, achieving an average improvement of 28.5% on CWEBench and a 140.8% improvement of Func-Sec@1 on CWEval. Results show that Secure-Instruct outperforms the SOTA method, demonstrating both scalability and efficiency.
Appendix A Appendix A: Prompt Template in Data Synthesizing
In our study, we leverage two schemes to synthesize data by vulnerable-code driven for Secure-Instruct(vul, secure) and secure-code driven for Secure-Instruct(vul, secure). In Secure-Instruct(vul, secure), Figure 5 shows the prompt template that we first prompt LLMs to generate vulnerable code snippets. Then another prompt is used for vulnerable code fixing, shown in Figure 6.
In Secure-Instructsecure, the prompt shown in Figure 7 is similar to the first step of Secure-Instruct(vul, secure), and the only change is that we directly ask LLMs to generate secure code snippets.
Prompt Template for Vulnerable Code Generation of Secure-Instruct(vul, secure) Task: Please generate an easily understandable vulnerable code snippet based on the following description of CWE. (Please consider the diversity of the vulnerable code examples for each generation.) CWE: {<CWE_OVERALL_DESCRIPTION>} A Vulnerable Code Example: This is {<LANGUAGE>} language example.
{<CODE>}
Figure 5: Prompt template used for vulnerable code generation in Secure-Instruct(vul, secure).
{<GENERATED_VULNERABLE_CODE>}
Can you FIX the code for {<LANGUAGE>} ? Please make sure the code is secure to the CWE mentioned above and runnable. Task: Please generate an easily understandable secure code snippet based on the following description of CWE. (Please consider the diversity of the vulnerable code examples for each generation.)
Figure 6: Prompt template used for vulnerable code fixing in Secure-Instruct(vul, secure).
Prompt Template for Secure Code Generation of Secure-Instructsecure Task: Please generate an easily understandable secure code snippet based on the following description of CWE. (Please consider the diversity of the secure code examples for each generation.) CWE: {<CWE_OVERALL_DESCRIPTION>} A Vulnerable Code Example: This is {<LANGUAGE>} language example.
{<CODE>}
Explanation of the Example: {<EXPLANATION>} Can you generate a secure code example for {<TARGET_LANGUAGE>} language?
Figure 7: Prompt template used for secure code generation in Secure-Instructsecure.
Prompt Template for Instruction Generation Create a single, very short (maximum two sentences) non-detailed functionality description that could be used as a prompt to generate either of the code snippets below… {<The_Non-vulnerable_code_snippet>}
Figure 8: Prompt template used for instruction generation.
Appendix B Appendix B: Distribution of CWEBench
Figure 9 shows the distribution of evaluated scenarios for languages and CWEs. Among 93 scenarios, 19 are for C, 9 are for Go, 14 are for Java, 18 are for JavaScript, 22 are for Python, and 11 are for Ruby. These scenarios cover 44 distinct CWEs, enabling a comprehensive evaluation.
Figure 9: The distribution of languages and CWEs in CWEBench.
Appendix C Appendix C: Detailed results of CWEBench
Table 6 breaks down the detailed results on CWEBench of CodeLlama-7B and Mistral-7B in their pre-trained, SafeCoder, and Secure-Instruct, on each CWE in our evaluation benchmark. The highest secure ratio for each CWE is highlighted in bold. In the scenarios whose CWEs are covered in both approaches (the top section of the table), our analysis shows that CodeLlama-7B fine-tuned with Secure-Instructsecure achieved the highest overall secure ratio (70.6%), outperforming SafeCoder by 2.3% in secure ratio, Mistral-7B fine-tuned with Secure-Instructsecure achieved 68.9% secure ratio, with a 3.7% higher than SafeCoder. In the generalization evaluation (i.e., with the full security-specific scenarios of our benchmark), Secure-Instructsecuredemonstrates strong generalization capabilities across a broader range of CWEs for both CodeLlama-7B and Mistral-7B. Specifically, CodeLlama-7B with Secure-Instructsecure significantly outperforms SafeCoder, achieving a 69.2% secure ratio, while Mistral-7B achieves a 66.4% secure ratio.
@@ Appendix C Appendix C: Detailed results of CWEBench
Pre-trained SafeCoder Secure-Instruct Secure Secure-Instruct (Vul, Secure) Pre-trained SafeCoder Secure-Instruct Secure Secure-Instruct (Vul, Secure)
@@
Appendix D Appendix D: Limitations
@@
LLMs can produce different outputs for the same prompt. To prevent the huge variations, we set a large (100) sample size to ensure the stabilization of results. CWEBench relies on CodeQL to detect vulnerabilities. Such a static analyzer may produce false positives. Although we manually reviewed samples, misclassifications could still influence the conclusions. We performed de-duplication across both our fine-tuning dataset and evaluation benchmarks (CWEBench and CWEval). However, this process may fail to detect duplicates, leading to false negatives.
@@
We follow prior work and choose LLMs with smaller parameter sizes to make fine-tuning feasible under limited computational resources. Instruction-tuning larger models, such as ChatGPT-4o, would incur significantly higher costs. As a result, we acknowledge that our findings may not fully generalize to larger LLMs, and we consider this a direction for future investigation. We include 44 distinct CWEs in this study, reflecting a broad range of vulnerabilities. However, numerous additional CWEs remain unaddressed in this study. Moreover, individual CWE can manifest in various real-world scenarios, further highlighting the complexity of comprehensive evaluation. To improve coverage, we expanded our security-sensitive prompts to incorporate additional CWEs beyond those included in earlier studies. Compared to prior work (pearce2022asleep; he2023large; he2024instruction; li2024exploratory), our study evaluates a significantly greater number of CWEs, providing a more thorough assessment of generated code vulnerabilities. Future efforts should include broader CWEs. Additionally, we include six programming languages in our evaluation benchmark. While these languages are highly popular and widely adopted in software development, our findings may not generalize to all programming languages. Future work should expand the language set to strengthen the robustness of the results.